OpenStudy (anonymous):

Is y(x-9) over x^2-9 already simplified? Can it not be simplified anymore? Answer me, please.

OpenStudy (lgbasallote):

remember that you can factor out x^2 - 9 further..do you know how to factor out difference of two squares?

OpenStudy (anonymous):

I don't know @lgbasallote

OpenStudy (lgbasallote):

Hint: \((a+b)(a-b) = a^2 - b^2\) that give you any ideas?

OpenStudy (anonymous):

Yeah

OpenStudy (lgbasallote):

good! so how can you factor out \(x^2 - 9^2\)

OpenStudy (anonymous):

The answer would be \[3y \over x+3\]. Am I right?

OpenStudy (lgbasallote):

hey wait...it's \(\frac {y(x-9)}{x^2 - 9}\)?

OpenStudy (anonymous):

Yeah

OpenStudy (lgbasallote):

well this is embarassing....i got ahead of myself..it cannot be simplified further..forgive me <tips imaginary hat>

OpenStudy (anonymous):

Ah. Ok. Thanks a lot

OpenStudy (lgbasallote):

sorry for wasting your time lol =)) this is sooo embarassing

OpenStudy (anonymous):

It's fine. It's good that I'm learning :) I actually thought it can still be simplified and came up with an answer. Haha